OpenStudy (anonymous):

Solve the system of equations using substitution. for a medal y = 3x + 1 y = 4x

OpenStudy (texaschic101):

y = 4x so sub 4x in for y in the other equation y = 3x + 1 4x = 3x + 1 -- subtract 3x from both sides 4x - 3x = 1 -- combine like terms x = 1 now sub 1 in for x in either equation to find y y = 4x y = 4(1) y = 4 now we can check our answers... y = 3x + 1 4 = 3(1) + 1 4 = 3 + 1 4 = 4 (correct) so x = 1 and y = 4 or (1,4) any questions ?